OpenStudy (anonymous):

how do I find 1/2 of 7 in fraction form?

OpenStudy (shane_b):

\[\frac{1}{2}*\frac{7}{1}=\frac{7}{2}\]

OpenStudy (ghazi):

\[\frac{ 1 }{ 2 }*7=\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

ok I thought so but it seemed so simple! T.Y

OpenStudy (ghazi):

:)

OpenStudy (shane_b):

It's very simple...sometimes you get lucky :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

the same should apply to 7/5 times 1/2?

OpenStudy (ghazi):

yes

OpenStudy (ghazi):

\[\frac{ 7 }{ 5 }*\frac{ 1 }{ 2 }=\frac{ 7 }{ 5*2 }\]

OpenStudy (anonymous):

okie dokie! ty again :)

OpenStudy (ghazi):

:) YW binky